OpenStudy (anonymous):

find the inverse laplace of 3 - 2s/(s^2 - 3s)^2

OpenStudy (anonymous):

F(s) = 1/(s(s - 3))

zepdrix (zepdrix):

What is this second post? Is that what it looks like simplified?

zepdrix (zepdrix):

Or just a separate problem?

OpenStudy (anonymous):

I found F'(s) = 3 - 2s/(S^2 - 3s)^2

OpenStudy (anonymous):

because f(t) = L^-1[F(s)] = -1/t * L^-1[F'(s)]

OpenStudy (anonymous):

s*

OpenStudy (anonymous):

The Problem: Find the inverse Laplace of F(s) = 1/(s(s-3)).

zepdrix (zepdrix):

Looks like we need to do Partial Fraction Decomposition. Remember that business? :O

OpenStudy (anonymous):

for which one? ugh

zepdrix (zepdrix):

\[\large \frac{1}{s(s-3)} \quad = \quad \frac{A}{s}+\frac{B}{s-3}\]Multiplying through by the denominator of the original gives us,\[\huge 1=A(s-3)+Bs\]Then you can solve for A and B from there :O Yes no? :D

OpenStudy (anonymous):

then once i have A and B?

OpenStudy (anonymous):

i can use the laplace table to determine inverse?

zepdrix (zepdrix):

Yah :) Through this process you'll be able to break up your fractions like this,\[\large \frac{A}{s}+\frac{B}{s-3}\]Where A and B are constants. And yes you should be able to find those on your table.

OpenStudy (anonymous):

okay sweet thanks